# High-Dimensional Gaussian Process Regression with Soft Kernel Interpolation

### v2

@@ 1 Introduction
Gaussian processes (GPs) are flexible function approximators based on Bayesian inference. However, there are scaling concerns. For a dataset comprising $n$ data points, constructing the $n\times n$ kernel (covariance) matrix required for GP inference incurs a space complexity of $\mathcal{O}(n^{2})$. Solving the associated linear system for exact posterior inference via direct methods requires $\mathcal{O}(n^{3})$ time complexity. These computational costs render exact GP inference challenging, even for modest $n$, although efforts have been made to accelerate exact GP inference (Wang et al., 2019; Gardner et al., 2021).
One approach to scalable GP regression is based in variational inference (Titsias, 2009; Hensman et al., 2013). These methods build a variational approximation of the posterior GP posterior by learning the locations of $m\ll n$ inducing points (Quinonero-Candela & Rasmussen, 2005; Snelson & Ghahramani, 2005). Inducing points and their corresponding inducing variables introduce latent variables with normal priors that form a low-rank approximation of the covariance structure in the original dataset. This approach improves the time complexity of posterior inference to $\mathcal{O}(m^{2}n)$ for Sparse Gaussian Process Regression (SGPR) (Titsias, 2009) and $\mathcal{O}(m^{3})$ for Stochastic Variational Gaussian Process Regression (SVGP) (Hensman et al., 2013), with the latter introducing additional variational parameters.
Another approach is based on Structured Kernel Interpolation (SKI) (Wilson & Nickisch, 2015) and its variants such as product kernel interpolation (SKIP) (Gardner et al., 2018) and Simplex-SKI (Kapoor et al., 2021). SKI-based methods achieve scalability by constructing computationally tractable approximate kernels via interpolation from a pre-computed and dense rectilinear grid of interpolation points. This structure enables fast matrix-vector multiplications (MVMs), which can be leveraged to scale inference via conjugate gradient (CG) methods. However, this approach also causes the time complexity of posterior inference to become explicitly dependent on the dimensionality $d$ of the data (e.g., $\mathcal{O}(n4^{d}+m\log m)$ for a MVM in SKI and $\mathcal{O}(d^{2}(n+m))$ for a MVM in Simplex-SKI). Moreover, the static grid does not have the flexibility of a variational method to adapt to the dataset at hand. This motivates us to search for accurate and scalable GP regression algorithms that can attain the best of both approaches.
@@
Our method approximates a kernel by interpolation from a softmax of $m\ll n$ learned interpolation points, hence the name soft kernel interpolation (Section 4.1). The interpolation points are learned by optimizing a combination of the SoftKI MLL, and when needed, an approximate MLL for improved numerical stability in single-precision floating point arithmetic (Section 4.2). It is able to leverage GPU acceleration and stochastic optimization for scalability. Since the kernel structure is dynamic, we rely on $m\ll n$ interpolation points to obtain a time complexity of $\mathcal{O}(m^{2}n)$ and space complexity of $\mathcal{O}(mn)$ for posterior inference (Section 4.3). We evaluate SoftKI on a variety of datasets from the UCI repository (Kelly et al., 2017) and demonstrate that it achieves test root mean square error (RMSE) comparable to other inducing point methods for datasets with moderate dimensionality (approximately $d=10$) (Section 5.1). To further explore the effectiveness of SoftKI in high dimensions, we apply SoftKI to molecule datasets ($d$ in the hundreds to thousands) and show that SoftKI is scalable and competitive in these settings as well (Section 5.2). Lastly, we compare the interpolation points learned by SoftKI to the inducing points learned by SGPR and SVGP, and find that they are structurally different, suggesting that learned interpolation points provide another promising path for constructing approximate GPs (Section 5.3).
2 Background
Notation. Matrices will be denoted by upper-case boldface letters (e.g. $\mathbf{A}\in\mathbb{R}^{n\times m}$). When specifying a matrix entry-wise we write $\mathbf{A}=[g(i,j)]_{i,j}$, meaning $\mathbf{A}_{ij}=g(i,j)$. At times, given points $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$, we stack them into $\mathbf{x}\in\mathbb{R}^{nd}$ by $\mathbf{x}^{\top}=(x_{1}^{\top},\dots,x_{n}^{\top}),$ so that entries $d(i-1)+1$ through $di$ are the components of $\mathbf{x}_{i}$. Any function $f:\mathbb{R}^{d}\to\mathbb{R}$ extends column-wise via $f(\mathbf{x})=(f(\mathbf{x}_{1}),\dots,f(\mathbf{x}_{n}))^{\top}.$
@@ 2.1 Gaussian Processes
Let $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ be a positive semi-definite kernel function. A (centered) Gaussian process (GP) with kernel $k$ is a distribution over functions $f:\mathbb{R}^{d}\to\mathbb{R}$ such that for any finite collection of inputs $\{x_{i}\}_{i=1}^{n}$, the vector of function values $f(\mathbf{x})$ is jointly Gaussian $f(\mathbf{x})\sim\mathcal{N}(\mathbf{0},\,\mathbf{K}_{\mathbf{x}\mathbf{x}})$, where the kernel matrix $\mathbf{K}_{\mathbf{x}\mathbf{x}}\in\mathbb{R}^{d\times d}$ has entries $[\mathbf{K}_{\mathbf{x}\mathbf{x}}]_{ij}=k(x_{i},\,x_{j}).$ Given two collections $\{x_{i}\}_{i=1}^{n}$ and $\{x^{\prime}_{j}\}_{j=1}^{n^{\prime}}$, the cross-covariance matrix $\mathbf{K}_{\mathbf{x}\mathbf{x}^{\prime}}\in\mathbb{R}^{d\times nd^{\prime}}$ is defined by $[\mathbf{K}_{\mathbf{x}\mathbf{x}^{\prime}}]_{ij}=k(x_{i},\,x^{\prime}_{j}).$
@@
$p(f(*)\,|\,\mathbf{y})=\mathcal{N}(\mathbf{K}_{*\mathbf{x}}(\mathbf{K}_{\mathbf{x}\mathbf{x}}+\mathbf{\Lambda})^{-1}\mathbf{y},\mathbf{K}_{**}-\mathbf{K}_{*\mathbf{x}}(\mathbf{K}_{\mathbf{x}\mathbf{x}}+\mathbf{\Lambda})^{-1}\mathbf{K}_{\mathbf{x}*})$ (1)
@@
[…]
@@ 2.2 Sparse Gaussian Process Regression
$\mathbf{K}_{\mathbf{x}\mathbf{x}}\approx\mathbf{K}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}^{-1}\mathbf{K}_{\mathbf{z}\mathbf{x}}$ (3)
leading to an overall inference complexity of $\mathcal{O}(m^{2}n)$.
@@
Variational inducing point methods such as SGPR (Titsias, 2009) combines the Nyström GP kernel approximation with a variational optimization process so that the positions of the inducing points can be learned. The variational approximation introduces latent inducing variables $u$ to model the values observed at inducing points $\mathbf{z}$ with joint distribution
$\begin{pmatrix}u\\ f(\mathbf{x})\end{pmatrix}$ $\sim\mathcal{N}\left(\begin{pmatrix}\mathbf{0}\\ \mathbf{0}\end{pmatrix},\begin{pmatrix}\mathbf{K}_{\mathbf{z}\mathbf{z}}&\mathbf{K}_{\mathbf{z}\mathbf{x}}\\ \mathbf{K}_{\mathbf{x}\mathbf{z}}&\mathbf{K}_{\mathbf{x}\mathbf{x}}\end{pmatrix}\right)\,.$ (4)
This leads to a posterior predictive distribution $q(f(*))$ which has closed form
$q(f(*)\,|\,\mathbf{y})=\mathcal{N}(f(*)\,|\,\mathbf{K}_{*\mathbf{z}}\mathbf{C}^{-1}\mathbf{K}_{\mathbf{z}\mathbf{x}}\mathbf{\Lambda^{-1}}\mathbf{y},\mathbf{K}_{**}-\mathbf{K}_{*\mathbf{z}}(\mathbf{K}_{\mathbf{z}\mathbf{z}}^{-1}-\mathbf{C^{-1}})\mathbf{K}_{\mathbf{z}*})\,.$ (5)
The time complexity of posterior inference is $\mathcal{O}(m^{2}n+m^{3})$ since it requires solving $\mathbf{C}\mathbf{\alpha}=\mathbf{K}_{\mathbf{z}\mathbf{x}}\mathbf{\Lambda}^{-1}\mathbf{y}$ for $\mathbf{\alpha}$ which is dominated by forming $\mathbf{C}$. The inducing points are learned by maximizing the evidence lower bound (ELBO)
@@
$\text{ELBO}(q)$ $=\log\mathcal{N}(\mathbf{y}\,|\,\mathbf{0},\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SGPR}}+\mathbf{\Lambda})-\frac{1}{2}\operatorname{tr}(\mathbf{K}_{\mathbf{x}\mathbf{x}}-\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SGPR}})$ (6)
@@
where $\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SGPR}}=\mathbf{K}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}^{-1}\mathbf{K}_{\mathbf{z}\mathbf{x}}$ is a Nyström approximation of $\mathbf{K}_{\mathbf{x}\mathbf{x}}$. Since the ELBO is a lower bound on the MLL $\log p(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$, we can maximize the ELBO via gradient-based optimization as a proxy for maximizing $\log p(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$ to learn the location of the inducing points $\mathbf{z}$ by treating it as a GP hyperparameter, i.e., $\mathbf{\theta}=(\beta,\ell,\sigma,\mathbf{z})$. The time complexity of computing the ELBO is $\mathcal{O}(m^{2}n)$.
@@ 2.3 Structured Kernel Interpolation
$\mathbf{K}_{\mathbf{x}\mathbf{x}^{\prime}}^{\text{SKI}}$ $=\mathbf{W}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}\mathbf{W}_{\mathbf{z}\mathbf{x}^{\prime}}\approx\mathbf{K}_{\mathbf{x}\mathbf{x}^{\prime}}$ (7)
@@
$\mathbf{K}_{\mathbf{x}\mathbf{x}^{\prime}}^{\text{SKI}}$ $=(\mathbf{W}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}})\mathbf{K}_{\mathbf{z}\mathbf{z}}^{-1}(\mathbf{K}_{\mathbf{z}\mathbf{z}}\mathbf{W}_{\mathbf{z}\mathbf{x}^{\prime}})=\mathbf{\widehat{K}}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}^{-1}\mathbf{\widehat{K}}_{\mathbf{z}\mathbf{x}^{\prime}}$ (8)
@@
$p(f(*)\,|\,\mathbf{y})\approx\mathcal{N}(\mathbf{K}_{*\mathbf{x}}(\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SKI}}+\mathbf{\Lambda})^{-1}\mathbf{y},\mathbf{K}_{**}-\mathbf{K}_{*\mathbf{x}}(\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SKI}}+\mathbf{\Lambda})^{-1}\mathbf{K}_{\mathbf{x}*})$ (9)
@@
where $\mathbf{K}_{\mathbf{x}\mathbf{x}}$ is replaced with $\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SKI}}$. SKI uses linear CG to solve the linear system $(\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SKI}}+\mathbf{\Lambda})\mathbf{\alpha}=\mathbf{y}$ for $\mathbf{\alpha}$. Since the time complexity of CG depends on access to fast MVMs, SKI makes two design choices that are compatible with CG. First, SKI uses cubic interpolation (Keys, 1981) weights so that $\mathbf{W}_{\mathbf{x}\mathbf{z}}$ has four non-zero entries in each row, i.e., $\mathbf{W}_{\mathbf{x}\mathbf{z}}$ is sparse. Second, SKI chooses the interpolation points $\mathbf{z}$ to be on a fixed rectilinear grid in $\mathbb{R}^{d}$ which induces multilevel Toeplitz structure in $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ if $k(x,y)$ is a stationary kernel.
@@
Putting the two together, the resulting $\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SKI}}$ is a structured kernel which enables fast MVMs. If we wish to have $k$ distinct interpolation points for each dimension, the grid will contain $m=k^{d}$ interpolation points leading to a total complexity of GP inference of $\mathcal{O}(n4^{d}+m\log(m))=\mathcal{O}(n4^{d}+dk^{d}\log(k))$ since there are $\mathcal{O}(4^{d})$ nonzero entries per row of $\mathbf{W}_{\mathbf{x}\mathbf{z}}$ and a MVM with a Toeplitz structured kernel like $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ can be done in $\mathcal{O}(m\log m)$ time via a Fast Fourier Transform (FFT) (Wilson, 2014). Thus, for low dimensional problems, SKI can provide significant speedups compared to vanilla GP inference and supports a larger number of interpolation points compared to SGPR.
3 Related Work
Beyond SGPR and vanilla SKI, a wide range of approximate GP models and variants have been developed that align closely with SoftKI.
Additional variational methods. Building on SGPR, SVGP (Hensman et al., 2013) extends the optimization process used for SGPR to use stochastic variational inference. This further reduces the computational cost of GP inference to $\mathcal{O}(m^{3})$ since optimization is now over a variational distribution on $\mathbf{z}$, rather than the full posterior. Importantly SVGP is highly scalable to GPUs since its optimization strategy is amenable to minibatch optimization procedures. More recent research has introduced Variational Nearest Neighbor Gaussian Processes (VNNGP) (Wu et al., 2024), which replace the low-rank prior of SVGP with a sparse approximation of the precision matrix by retaining only correlations among each point’s $K$ nearest neighbors. By constructing a sparse Cholesky factor with at most $K+1$ nonzeros per row, VNNGP can reduce the per-iteration cost of evaluating the SVGP objective Equation 6 to $\mathcal{O}\bigl{(}(n_{b}+m_{b})K^{3}\bigr{)}$ when minibatching over $n_{b}$ data points and $m_{b}$ inducing points.
[…]

### v3

@@ 1 Introduction
Gaussian processes (GPs) are flexible function approximators based on Bayesian inference. However, there are scaling concerns. For a dataset comprising $n$ data points, constructing the $n\times n$ kernel (covariance) matrix required for GP inference incurs a space complexity of $\mathcal{O}(n^{2})$. Naively, posterior inference using direct solvers costs $\mathcal{O}(n^{3})$ time, which quickly becomes infeasible for moderate $n$. In recent years, this shortcoming has led to the widespread adoption of conjugate gradient (CG) based GPs enabling “exact" GP inference in $\mathcal{O}(n^{2})$ time (Cunningham et al., 2008; Cutajar et al., 2016). Despite these developments, exact GP regression remains challenging in practice, often requiring bespoke engineering solutions (Wang et al., 2019; Gardner et al., 2018a).
A popular response to this scaling issue is based on variational inference methods (Titsias, 2009; Hensman et al., 2013). These approaches build a variational approximation of the posterior GP by learning the locations of $m\ll n$ inducing points (Quinonero-Candela & Rasmussen, 2005; Snelson & Ghahramani, 2005). Inducing points and their corresponding inducing variables introduce latent variables with normal priors that form a low-rank approximation of the covariance structure in the original dataset. This approach improves the time complexity of posterior inference to $\mathcal{O}(nm^{2})$ for Sparse Gaussian Process Regression (SGPR) (Titsias, 2009) and $\mathcal{O}(m^{3})$ for Stochastic Variational Gaussian Process Regression (SVGP) (Hensman et al., 2013), with the latter introducing additional variational parameters.
Another successful approach is based on Structured Kernel Interpolation (SKI) (Wilson & Nickisch, 2015) and its variants such as product kernel interpolation (SKIP) (Gardner et al., 2018b), Simplex-SKI (Kapoor et al., 2021), and Sparse-Grid SKI (Yadav et al., 2022). SKI-based methods achieve scalability by constructing computationally tractable approximate kernels via interpolation from a pre-computed and dense rectilinear grid of interpolation points. This structure enables fast matrix-vector multiplications (MVMs), which leads to downstream acceleration when paired with CG-based kernel inversion strategies. Superficially, this approach effectively responds to the aforementioned scaling issues, but unfortunately has the consequence of causing the complexity of posterior inference to become explicitly dependent on the dimensionality $d$ of the data (e.g., $\mathcal{O}(n4^{d}+dm^{d}\log m)$ for a MVM in SKI and $\mathcal{O}(d^{2}(n+m))$ for a MVM in Simplex-SKI). Moreover, the static grid used in SKI approximations does not have the flexibility of a variational method to adapt to the dataset at hand. This limitation motivates us to search for accurate and scalable GP regression algorithms that can attain the best of both approaches.
@@
Our method approximates a kernel by interpolation from a softmax of $m\ll n$ learned interpolation points, hence the name soft kernel interpolation (Section 3.1). The interpolation points are learned by optimizing a combination of the SoftKI MLL, and when needed, an approximate MLL for improved numerical stability in single-precision floating point arithmetic (Section 3.2). It is able to leverage GPU acceleration and stochastic optimization for scalability. Since the kernel structure is dynamic, we rely on $m\ll n$ interpolation points to obtain a time complexity of $\mathcal{O}(nm^{2})$ and space complexity of $\mathcal{O}(nm)$ for posterior inference (Section 3.3). We evaluate SoftKI on a variety of datasets from the UCI repository (Kelly et al., 2017) and demonstrate that it achieves test root mean square error (RMSE) comparable to other inducing point methods for datasets with moderate dimensionality (approximately $d=10$) (Section 4.1). To further explore the effectiveness of SoftKI in high dimensions, we apply SoftKI to molecule datasets ($d$ in the hundreds to thousands) and show that SoftKI is scalable and competitive in these settings as well (Section 4.2). Lastly, we examine the numerical stability of SoftKI (Section 4.3).
2 Background and Related Work
Notation. Matrices will be denoted by upper-case boldface letters (e.g. $\mathbf{A}\in\mathbb{R}^{n\times m}$). When specifying a matrix entry-wise generated by an underlying function we write $\mathbf{A}=[g(i,j)]_{i,j}$, meaning $\mathbf{A}_{ij}=g(i,j)$. At times, given points $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$, we stack them into $\mathbf{x}\in\mathbb{R}^{nd}$ by $\mathbf{x}^{\top}=(x_{1}^{\top},\dots,x_{n}^{\top}),$ so that entries d(i1)+1d(i-1)+1 through $di$ are the components of $\mathbf{x}_{i}$. Any function $f:\mathbb{R}^{d}\to\mathbb{R}$ extends column-wise via $f(\mathbf{x})=(f(\mathbf{x}_{1}),\dots,f(\mathbf{x}_{n}))^{\top}.$
@@ 2.1 Gaussian Processes
Let $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ be a positive semi-definite kernel function. A (centered) Gaussian process (GP) with kernel $k$ is a distribution over functions $f:\mathbb{R}^{d}\to\mathbb{R}$ such that for any finite collection of inputs $\{x_{i}\}_{i=1}^{n}$, the vector of function values $f(\mathbf{x})$ is jointly Gaussian $f(\mathbf{x})\sim\mathcal{N}(\mathbf{0},\,\mathbf{K}_{\mathbf{x}\mathbf{x}})$, where the kernel matrix $\mathbf{K}_{\mathbf{x}\mathbf{x}}\in\mathbb{R}^{n\times n}$ has entries $\mathbf{K}_{\mathbf{x}\mathbf{x}}]_{ij}=k(x_{i},\,x_{j}).$ Given two collections $\{x_{i}\}_{i=1}^{n}$ and $\{x^{\prime}_{j}\}_{j=1}^{n^{\prime}}$, the cross-covariance matrix is denoted $\mathbf{K}_{\mathbf{x}\mathbf{x}^{\prime}}\in\mathbb{R}^{n\times n^{\prime}}$ is defined by $\mathbf{K}_{\mathbf{x}\mathbf{x}^{\prime}}]_{ij}=k(x_{i},\,x^{\prime}_{j}).$
@@
$p(f(*)\,|\,\mathbf{x},\mathbf{y})=\mathcal{N}(\mathbf{K}_{*\mathbf{x}}(\mathbf{K}_{\mathbf{x}\mathbf{x}}+\mathbf{\Lambda})^{-1}\mathbf{y},\mathbf{K}_{**}-\mathbf{K}_{*\mathbf{x}}(\mathbf{K}_{\mathbf{x}\mathbf{x}}+\mathbf{\Lambda})^{-1}\mathbf{K}_{\mathbf{x}*})$ (1)
@@
$\log p(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})=\log\mathcal{N}(\mathbf{y}\,|\,\mathbf{0},\mathbf{K}_{\mathbf{x}\mathbf{x}}(\mathbf{\theta})+\mathbf{\Lambda}(\mathbf{\theta}))$
@@ 2.2 Sparse Gaussian Process Regression
$\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SGPR}}=\mathbf{K}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}^{-1}\mathbf{K}_{\mathbf{z}\mathbf{x}}\approx\mathbf{K}_{\mathbf{x}\mathbf{x}},$
leading to an overall inference complexity of $\mathcal{O}(nm^{2})$.
@@
Variational inducing point methods such as SGPR (Titsias, 2009) combines the Nyström GP kernel approximation with a variational optimization procedure so that the positions of the inducing points can be learned. The variational approximation introduces latent inducing variables $f(\mathbf{z})$ to model the values observed at inducing points $\mathbf{z}$ with joint distribution
$p(f(\mathbf{x}),f(\mathbf{z}))=\mathcal{N}\!\left(\begin{pmatrix}f(\mathbf{x})\\ f(\mathbf{z})\end{pmatrix}\,\middle|\,\begin{pmatrix}\mathbf{0}\\ \mathbf{0}\end{pmatrix},\begin{pmatrix}\mathbf{K}_{\mathbf{x}\mathbf{x}}&\mathbf{K}_{\mathbf{x}\mathbf{z}}\\ \mathbf{K}_{\mathbf{z}\mathbf{x}}&\mathbf{K}_{\mathbf{z}\mathbf{z}}\end{pmatrix}\right)$.
The inducing points are learned by maximizing the evidence lower bound (ELBO)
$\operatorname{ELBO}(q)=\mathbb{E}_{q(\mathbf{Z})}\bigl{[}\log(p(\mathbf{Y}\,|\,\mathbf{Z})\bigr{]}-\operatorname{KL}(q(\mathbf{Z})\,||\,p(\mathbf{Z}))$ (2)
where $\operatorname{KL}$ is the KL divergence between two probability distributions, $q$ is a computationally tractable variational distribution, $\mathbf{Y}$ is observed data ($\mathbf{Y}=\mathbf{y}$ for SGPR), and $\mathbf{Z}$ are latent variables ($\mathbf{Z}=(f(\mathbf{x}),f(\mathbf{z}))$ for SGPR). In the case of SGPR regression, the variational family $q(f(\mathbf{x}),f(\mathbf{z}))=p(f(\mathbf{x})\,|\,f(\mathbf{z}))q(f(\mathbf{z}))$ is chosen to approximate $p(f(\mathbf{x}),f(\mathbf{z}))$ so that Equation 2 simplifies to
@@
$\operatorname{ELBO}(q)=\log\mathcal{N}\left(\mathbf{y}\mid\mathbf{0},\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SGPR}}+\mathbf{\Lambda}\right)-\frac{1}{2}\operatorname{tr}\left(\mathbf{K}_{\mathbf{x}\mathbf{x}}-\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SGPR}}\right).$
@@
Since the ELBO is a lower bound on the MLL $\log p(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$, we can maximize the ELBO via gradient-based optimization as a proxy for maximizing $\log p(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$ to learn the location of the inducing points $\mathbf{z}$ by treating it as a GP hyperparameter, i.e., $\mathbf{\theta}=(\beta,\ell,\sigma,\mathbf{z})$. The time complexity of computing the ELBO is $\mathcal{O}(nm^{2})$.
The posterior predictive distribution has closed form
$q(f(*)\,|\,\mathbf{x},\mathbf{y})=\mathcal{N}(f(*)\,|\,\mathbf{K}_{*\mathbf{z}}\mathbf{C}^{-1}\mathbf{K}_{\mathbf{z}\mathbf{x}}\mathbf{\Lambda^{-1}}\mathbf{y},\mathbf{K}_{**}-\mathbf{K}_{*\mathbf{z}}(\mathbf{K}_{\mathbf{z}\mathbf{z}}^{-1}-\mathbf{C^{-1}})\mathbf{K}_{\mathbf{z}*})$
where $\mathbf{C}=\mathbf{K}_{\mathbf{z}\mathbf{z}}+\mathbf{K}_{\mathbf{z}\mathbf{x}}\mathbf{\Lambda^{-1}}\mathbf{K_{\mathbf{x}\mathbf{z}}}$. The time complexity of posterior inference is $\mathcal{O}(nm^{2}+m^{3})$ since it requires solving $\mathbf{C}\mathbf{\alpha}=\mathbf{K}_{\mathbf{z}\mathbf{x}}\mathbf{\Lambda}^{-1}\mathbf{y}$ for $\mathbf{\alpha}$, which is dominated by the cost of forming $\mathbf{C}$.
Additional variational methods. Building on SGPR, SVGP (Hensman et al., 2013) extends the optimization process used for SGPR to use stochastic variational inference. This further reduces the computational cost of GP inference to $\mathcal{O}(m^{3})$ since optimization is now over a variational distribution on $\mathbf{z}$, rather than the full posterior. Importantly SVGP is highly scalable to GPUs since its optimization strategy is amenable to minibatch optimization procedures. More recent research has introduced Variational Nearest Neighbor Gaussian Processes (VNNGP) (Wu et al., 2022), which replaces the low-rank prior of SVGP with a sparse approximation of the precision matrix by retaining only correlations among each point’s $K$ nearest neighbors. By constructing a sparse Cholesky factor with at most $K+1$ nonzeros per row, VNNGP can reduce the per-iteration cost of evaluating the SVGP objective (Equation 2.2) to $\mathcal{O}\bigl{(}(n_{b}+m_{b})K^{3}\bigr{)}$ when minibatching over $n_{b}$ data points and $m_{b}$ inducing points.
@@ 2.3 Structured Kernel Interpolation
$\mathbf{K}_{\mathbf{x}\mathbf{x}^{\prime}}^{\text{SKI}}=\mathbf{W}_{\mathbf{x}\mathbf{z}}\,\mathbf{K}_{\mathbf{z}\mathbf{z}}\,\mathbf{W}_{\mathbf{z}\mathbf{x}}\approx\mathbf{K}_{\mathbf{x}\mathbf{x}}.$
@@
$\mathbf{K}_{\mathbf{x}\mathbf{x}^{\prime}}^{\text{SKI}}=\left(\mathbf{W}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}\right)\mathbf{K}_{\mathbf{z}\mathbf{z}}^{-1}\left(\mathbf{K}_{\mathbf{z}\mathbf{z}}\mathbf{W}_{\mathbf{z}\mathbf{x}}\right)=\widehat{\mathbf{K}}_{\mathbf{x}\mathbf{z}}\,\mathbf{K}_{\mathbf{z}\mathbf{z}}^{-1}\,\widehat{\mathbf{K}}_{\mathbf{z}\mathbf{x}}.$
@@
$p(f(*)\,|\,\mathbf{x},\mathbf{y})\approx\mathcal{N}(\mathbf{K}_{*\mathbf{x}}(\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SKI}}+\mathbf{\Lambda})^{-1}\mathbf{y},\mathbf{K}_{**}-\mathbf{K}_{*\mathbf{x}}(\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SKI}}+\mathbf{\Lambda})^{-1}\mathbf{K}_{\mathbf{x}*})$ (3)
@@
where $\mathbf{K}_{\mathbf{x}\mathbf{x}}$ is replaced with $\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SKI}}$. When Equation 3 is solved via linear CG with structured interpolation (e.g., cubic), the resulting method is called KISS-GP (Wilson & Nickisch, 2015). Since the time complexity of CG depends on access to fast MVMs, KISS-GP makes two design choices that are compatible with CG. First, KISS-GP uses cubic interpolation weights (Keys, 1981) so that $\mathbf{W}_{\mathbf{x}\mathbf{z}}$ has only four non-zero entries in each row, i.e., $\mathbf{W}_{\mathbf{x}\mathbf{z}}$ is sparse. Second, KISS-GP chooses the interpolation points $\mathbf{z}$ to be on a fixed rectilinear grid in $\mathbb{R}^{d}$ which induces multilevel Toeplitz structure in $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ if $k(x,y)$ is a stationary kernel.
@@
Putting the two together, the resulting kernel $\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SKI}}$ is a highly structured structured matrix which enables fast MVMs. If we wish to have $m$ distinct interpolation points for each dimension, the grid will contain $m^{d}$ interpolation points leading to a total complexity of GP inference of $\mathcal{O}(n4^{d}+dm^{d}\log(m))$ since there are $\mathcal{O}(4^{d})$ nonzero entries per row of $\mathbf{W}_{\mathbf{x}\mathbf{z}}$ and a MVM with a Toeplitz structured kernel like $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ can be done in $\mathcal{O}(|\mathbf{z}|\log(|\mathbf{z}|))$ time via a Fast Fourier Transform (FFT) (Wilson, 2014). Thus, for low-dimensional problems, SKI can provide significant speedups compared to vanilla GP inference and supports a larger number of interpolation points compared to SGPR.
Table 1: Cost of a single MVM with the approximate covariance for various GP approximations. For listed methods based on SKI we take $m$ points in each dimension. The variable $r$ represents the rank used in the SKIP approximation (typically $r\approx 100$), and the variable $\ell$ is the grid resolution taken in Sparse-Grid GP (typically $\ell\approx 1-5$).
Method Time Complexity of one MVM
KISS-GP (cubic) $\mathcal{O}\bigl{(}n4^{d}+dm^{d}\log(m)\bigr{)}$
SKIP $\mathcal{O}\bigl{(}d(rn+m\log m\bigr{)})$
Simplex-GP $\mathcal{O}\bigl{(}d^{2}(n+m))$
Sparse-Grid GP (simplicial) $\mathcal{O}\bigl{(}nd^{2+\ell}+\ell^{d}2^{\ell})$
SGPR $\mathcal{O}\bigl{(}nm^{2}+m^{3}\bigr{)}$
SoftKI $\mathcal{O}\bigl{(}nm^{2}+m^{3}\bigr{)}$
2.4 Other Variants of Structured Kernel Interpolation.
While SKI can provide significant speedups over vanilla GP inference, its scaling in $d$ restricts its usage to low-dimensional settings where $4^{d}<n$. To address this, Gardner et al. (2018b) introduced SKIP, which optimizes SKI by expressing a $d$-dimensional kernel as a product of $d$ one-dimensional kernels. This reduces the MVM cost with $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ from $\mathcal{O}(dm^{d}\log(m))$ to $\mathcal{O}(dm\log(m))$, using only $m$ grid points per component kernel instead of $m^{d}$. However, SKIP is limited to dimensions roughly in the range of $d=10\text{--}30$ for large datasets, as it requires substantial memory and may suffer from low-rank approximation errors.